# Integral modelling and Reinforcement Learning control of 3D liquid metal coating on a moving substrate

### v1

@@ VII.1 Asymptotic Behaviour
In this subsection, we analyse the behaviour of the steady-state solution found in Subsection III.2 and the closure relations introduced in Subsection III.4 as the magnetic field $\hat{b}$ and the Hartmann number Ha approach zero.
@@
The velocity reaches a well-defined limiting value, as it recovers, at leading order, the solution that was previously derived in the case without magnetic field by Ivanova et al. [34]. This result serves as the first analytical validation of our model, confirming its asymptotic consistency.
Figure 9 presents the steady-state results for case (a) without and (b) with a magnetic field ($\mbox{Ha}=3$), for different values of $\hat{h}$. In both cases, the velocity profile is parabolic, but the variation between the different profiles is more pronounced in the absence of the magnetic field. In particular, the magnetic field has a more significant effect when $\hat{h}>1$, where the streamwise velocity remains negative, balancing the gravitational effects. In contrast, without the magnetic field, the downward gravitation force causes the velocity to become positive for most of the liquid film thickness.
@@
Moving to the steady-state flow rate (18), Figure 10 shows the relationship between $\bar{q}$ and $\bar{h}$ for different values of Ha (lines with markers) and the case without a magnetic field (black continuous line). Like the velocity profile, $\bar{q}$ recovers the case without a magnetic field as $\mbox{Ha}\to 0$. In the absence of the magnetic field, $\bar{q}$ has two solution branches: thin film ($\hat{h}<1$) and thick film ($\sqrt{3}\geq\hat{h}>1$), with a minimum at $\bar{h}=1$, corresponding to the flat film solution of Derjaguin [53]. The branch with $\bar{q}>0$ in the range $\hat{h}>\sqrt{3}$ is attainable solely by withdrawing the substrate from the bath and, at the same time, feeding the liquid film from above.
When the magnetic field is applied, the minimum moves towards smaller flow rates and thinner film thicknesses as Ha increases. In addition, the position of the zero net flow rate changes to higher values of $\bar{h}$ as Ha increases. This indicates that the locus of possible steady-state solutions, attainable by simply extracting the liquid film from a bath, expands with the application of the magnetic field.
@@
$\begin{split}F_{12}=&\frac{\hat{h}^{4}\hat{\tau}_{gx}^{2}+24\left(\hat{h}^{2}+2\hat{h}\hat{q}_{x}+6\hat{q}_{x}^{2}\right)}{120\hat{h}}\\ &+\frac{6\hat{h}^{2}\hat{\tau}_{gx}(\hat{h}+\hat{q}_{x})}{120\hat{h}}+O\left(\mbox{Ha}^{2}\right)\hat{b}+O\left(\hat{b}^{2}\right)\end{split}$ (52a)
$\begin{split}F_{13}=&\frac{\hat{h}^{4}\hat{\tau}_{gx}\hat{\tau}_{gz}+3\hat{h}^{2}\hat{\tau}_{gz}(\hat{h}+\hat{q}_{x})}{120\hat{h}}\\ &+\frac{3\hat{h}\hat{q}_{z}(\hat{h}\hat{\tau}_{gx}+8)+144\hat{q}_{x}\hat{q}_{z}}{120\hat{h}}\\ &+O\left(\mbox{Ha}^{2}\right)\hat{b}+O\left(\hat{b}^{2}\right)\,,\end{split}$ (52b)
$\begin{split}F_{23}=&\frac{\hat{h}^{4}\hat{\tau}_{gz}^{2}+6\hat{h}^{2}\hat{q}_{z}\hat{\tau}_{gz}+144\hat{q}_{z}^{2}}{120\hat{h}}\\ &-O\left(\mbox{Ha}^{2}\right)\hat{b}+O(\hat{b}^{2})\,,\end{split}$ (52c)
$\begin{split}\hat{\tau}_{w,x}=&\left(\frac{\hat{\tau}_{g,x}}{2}-\frac{3(\hat{h}+\hat{q}_{x})}{\hat{h}^{2}}\right)+O(\mbox{Ha}^{2})\hat{b}+O(\hat{b}^{2})\,,\end{split}$ (52d)
$\begin{split}\hat{\tau}_{w,z}=&\left(\frac{\hat{\tau}_{g,x}}{2}-\frac{3\hat{q}_{z}}{\hat{h}^{2}}\right)+O(\mbox{Ha}^{2})\hat{b}+O(\hat{b}^{2})\,.\end{split}$ (52e)
@@ VII.2 Leading-Oder convective velocity
As for the cases treated in subsection VII.1, the leading-order term is consistent with the case without magnetic field [34].
@@
Figure 11 illustrates the phase speed $c$ (54) as a function of the non-dimensional thickness $\hat{h}$ and the Hartmann number Ha. For small values of Ha, $c$ has a quadratic dependence on $\hat{h}$, with waves propagating upstream ($c<0$) for $\hat{h}<1$ and downstream ($c>0$) for $\hat{h}>1$. For $\mbox{Ha}\geq 2$, $c$ reaches a plateau at approximately $c\approx 0.1$ for $\hat{h}>0.3$. As we have already seen in figure 9(b), the magnetic effects for these conditions counterbalance the gravitational effects, resulting in an equilibrium solution where waves always propagate in the direction of substrate movement. This highlights how the magnetic force might affect the absolute/convective instability properties of these flat film solutions compared to the case without a magnetic field [54].
@@ VII.3 Undulation control using Reinforcement Learning
In this section, we present the results of the undulation control problem outlined in subsection VI.2, using the PPO algorithm introduced in subsection VI.1. We examine the cases with a single gas jet or electromagnetic actuators and the cases where these are used together. The results are presented in terms of learning curves, the mean, and standard deviation over the testing episodes, the evolution of the control function over time, and snapshots of the liquid film along the centreline.
VII.3.1 Control with 2D gas-jets
The first case involves a single 2D gas jet actuator in the case of a prescribed harmonic control function and in the case where the agent is free to take any action in time.
Figure 12: Learning curves for the harmonic (green line with squares) and non-harmonic (blue line with circles) control functions with a 2D gas jet.
@@ VII.3.1 Control with 2D gas-jets
Table 2 summarises the mean and standard deviation of the evaluation episodes for controlled and uncontrolled cases. Like in the learning phase, both control functions result in a higher mean reward than in the uncontrolled case. Furthermore, in the controlled case, the reward variability around the mean is reduced, with a smaller standard deviation. This shows how the PPO managed to find an effective control action that led to a reduced undulation amplitude in the reward area.
Focussing on the harmonic case, the evolution of the optimal control function during an evaluation episode is reported in figure 13. Figure 13(a) shows the amplitude (solid red line with squares), the nondimensional frequency (blue dashed line with triangles), and the phase shift (dotted dashed line with circles) of the harmonic control function during an evaluation episode. These three parameters remain nearly constant, with an amplitude of $\acute{A}\approx 0.68$, a frequency of $\acute{f}\approx 0.065$, and a phase shift of $\acute{\phi}\approx\pi$. The frequency aligns with the average frequency of the undulation waves, which ranges between $0.05$ and $0.08$. This shows how a single harmonic function is sufficient to reduce the amplitude of the undulation instability.
@@
To gain a better understanding of the control mechanism, we look at the evolution of the control function compared to the observation of the liquid film collected at one of the three observation locations. Figure 13(b) shows the evolution of the nozzle exit velocity (solid red line) and the observation at obs2 (solid green line). The positions of the crests and valleys of the two functions are very similar, with only a small phase shift because the observation point is located ahead of the control actuator. This shows that the jet primarily flattens the crests with only a minor effect on the valleys. However, since the nozzle exit velocity is always non-zero, this indicates that the optimal control law involves a thinning of the liquid film and a harmonic component nearly in phase with the instability wave.
Figure 15: Evolution of (a) the controlled 3D liquid film (a, b and c) and (b) along $x$ at $z=1[\rm{cm}]$ in the controlled (continuous blue line) and the uncontrolled (dashed green line) cases with a single electromagnetic actuator.
Moving to the case where the PPO is free to take any actions, figure 14 shows the evolution of (a) the control actions and the relation between the control actions and (b) the obs1 with the fitted line (red curve) for the non-harmonic control function. The control action is proportional to the mean liquid film thickness at the impingement point with a small phase shift.
@@
VII.3.2 Control with 2D electromagnets
Turning to the control with a single 2D electromagnetic actuator, Figure 16 shows the learning curve (solid black line), the mean value (red dashed line) and the standard deviation (yellow shaded area) of the reward in the absence of control. The case with a prescribed harmonic control function fails to produce good results, with the average reward well below that of the uncontrolled case, even falling outside the uncertainty region of the no-control case. This suggests that the agent could not find an efficient control law and continued to test random combinations of actions, which led to an increased amplitude of undulation within the reward area. In contrast, the agent discovered an optimal non-harmonic control law in the case without a prescribed control function shape at the 40th learning episode. This episode marks the point at which the mean reward surpassed the mean reward of the uncontrolled case. Furthermore, the mean value closely approaches the upper bound of the uncertainty region for the uncontrolled case, underscoring the robust performance achieved by the algorithm.
@@ VII.3.2 Control with 2D electromagnets
A similar behaviour can be observed when examining the performance during the evaluation episodes. Table 3 shows the mean and standard deviation of the liquid film with and without control, using both harmonic and non-harmonic control functions with the 2D electromagnetic actuator. The harmonic control function resulted in a mean reward that was 14% lower than the uncontrolled case, with a larger standard deviation. In contrast, the non-harmonic control function yielded a reward 8% higher than the uncontrolled case, with a smaller standard deviation.
@@
The control strategy obtained with the electromagnets differs significantly from the one we found using a single 2D gas jet. Figure 17 shows (a) the evolution of the control function (solid red line with squares) and the observations (coloured lines without markers) and (b) the relationship between the control function and obs1 during an evaluation episode. The electromagnetic control exhibits a bang-bang-type behaviour, acting only near the valleys with a maximum intensity of $0.14\,\rm{T}$. The actions are inversely proportional to obs1 when they exceed the flat state of the target, $\bar{h}=0.1\,(34.2\,\mu\rm{m})$.
Figure 17: Evolution of the observation and actions (red continuous curve with squares) (a) and the relation between the actions and the second observations (b) of the 3D undulation control with 2D electromagnets.
@@
The optimal control law mechanism is due to the resistive effects of the Lorentz force on the liquid film. The Lorenz forces push the liquid in the positive $x$ direction, which creates a positive/negative flow rate gradient upstream/downstream of the electromagnet centerline. This leads to an increase/decrease in the liquid film thickness upstream and a decrease downstream. This behaviour is evident when comparing the evolution of the waves in the controlled and uncontrolled cases. Figure 18 shows the plot along $x$ for the liquid film with (solid blue line) and without (dashed green line) control. The action of the electromagnet raises and deforms the valleys, which reduces the amplitude of the nearby peaks.
This control mechanism is limited by the rapid dynamics of the liquid film, which does not provide sufficient time for the Lorentz force to flatten the free surface. As shown in Table 1, the Stuart number N, which represents the relative importance of the Lorentz force compared to inertia, is well below one ($\mbox{N}=1.9\times 10^{-2}$). The squared ratio between the electromagnetic and inertial time scales is given by the product of N by the film parameter $\varepsilon$ as presented at the end of Section III.1. In our case, we obtain that the characteristic inertia timescale $t_{\rm ref}$ is three times smaller than the Lorentz force timescale $t_{\rm ref,M}$:
$t_{\rm ref}=\frac{\sqrt{N}}{\mbox{Ca}^{1/6}}t_{\rm ref,M}=0.37t_{\rm ref,M}$ (56)
For more efficient flattening of the liquid film, the magnetic timescale should be of the same order of magnitude as the inertial timescale, which would be possible to impose a Hartmann number given by the following relationship:
$\mbox{Ha}=\sqrt{\mbox{Re}}\mbox{Ca}^{1/6}=16.$ (57)
@@
Figure 18: Evolution of the controlled (blue continuous line) and the uncontrolled (green dashed line) along $x$ at $z=1[\rm{cm}]$ for the control of the 3D undulation with a 2D electromagnetic actuator.
VII.3.3 Control with 2D gas jets and electromagnets
@@
In the previous test cases, we found that optimal control actions of the gas jet and electromagnetic actuators complement each other. The gas jet pushes the crests, whereas the electromagnets raise the valleys. In this case, we test the control of undulations with both actuators working together. Additionally, to simplify the control problem, the agent has access to only three liquid film thickness observations instead of six.
@@
Figure 19 shows the learning curve along with the mean (red dashed line) and standard deviation (yellow shaded area) of the reward obtained without control. The PPO algorithm finds an optimal control law that outperforms the uncontrolled case after 100 training episodes. Furthermore, as observed in the case with a single electromagnet, the learning curve approaches the upper bound of the uncertainty region, demonstrating the robustness of the discovered control action. Similar results are evident in terms of evaluation episodes.
@@
Table 4 shows the mean and standard deviation of the reward obtained during the evaluation episodes, both with and without control. The optimal control law yields a mean reward 13% higher than in the uncontrolled case, and a 20% reduction in the standard deviation. This indicates that even in the worst evaluation test case, the agent outperforms the uncontrolled case on average.
Moving to the evolution of optimal control functions during an evaluation episode, Figure 20 shows (a) the actions of the 2D gas jet (blue solid line) and obs3 (green dashed line) and (b) the actions of the 2D electromagnet (red solid line) and obs2 (green dashed line). Similarly to the case with a single gas jet, the optimal control law is nearly proportional to the liquid film observation, with a small phase shift owing to the distance between the observation point and the actuator. In contrast, the magnetic control action differs significantly from that in the standalone case. Here, the magnetic field is never zero and changes in a non-proportional way compared to the observation, unlike the gas jet’s upstream action, which alters the free surface topography and forces the electromagnet to adopt a different control strategy. Additionally, since both actuators are always nonzero, the optimal control action implies further thinning of the liquid film.
The control mechanism, which combines pushing crests and raising valleys, is most apparent when examining the evolution of the liquid film along a line in the streamwise direction. Figure 21 shows the evolution of the liquid film with control (solid blue line) and without control (dashed green line) along $x$ at $z=1\,[\rm{cm}]$. For a single actuator, the gas jets push down the crests with additional wiping. However, this is counteracted by the action of the electromagnet, resulting in waves with a smaller peak-to-valley amplitude in the controlled case compared to the uncontrolled one.

### v2

@@ VII.1 Asymptotic Behaviour
This subsection focuses on the behaviour of the steady-state solution found in Subsection III.2 and the closure relations introduced in Subsection III.4 as the magnetic field $\hat{b}$ and the Hartmann number Ha approach zero.
@@
The velocity reaches a well-defined limiting value, as it recovers, at leading order, the solution that was previously derived in the case without a magnetic field by Ivanova et al. [22]. This result serves as the first analytical validation of our model, confirming its asymptotic consistency.
Figure 6 presents the steady-state results for case (a) without and (b) with a magnetic field ($\mbox{Ha}=3$), for different values of $\hat{h}$. In both cases, the velocity profile is parabolic, but the variation between the different profiles is more pronounced in the absence of the magnetic field. In particular, the magnetic field exerts a stronger effect when $\hat{h}>1$, in which case the streamwise velocity remains negative, balancing the gravitational force. In contrast, in the absence of the magnetic field, the downward gravitational force causes the velocity to become positive over most of the liquid film thickness.
@@
Moving to the steady-state flow rate (22), Figure 7 shows the relationship between $\bar{q}$ and $\bar{h}$ for different values of Ha (lines with markers) and the case without a magnetic field (black continuous line). Like the velocity profile, $\bar{q}$ recovers the case without a magnetic field as $\mbox{Ha}\to 0$. In the absence of the magnetic field, $\bar{q}$ is equal to zero for $\hat{h}=0$ and $\hat{h}=\sqrt{3}$ and it has two solution branches: thin film ($\hat{h}<1$) and thick film ($\sqrt{3}\geq\hat{h}>1$), with a minimum at $\bar{h}=1$, corresponding to the flat film solution of Derjaguin [12]. The branch with $\bar{q}>0$ in the range $\hat{h}>\sqrt{3}$ is attainable solely by withdrawing the substrate from the bath and, at the same time, feeding the liquid film from above.
When the magnetic field is applied, the minimum shifts to lower flow rates and thinner film thicknesses as Ha increases. In addition, the position of the zero net flow rate changes to higher values of $\bar{h}$ as Ha increases. This indicates that the locus of possible steady-state solutions, attainable by simply extracting the liquid film from a bath, expands with the application of the magnetic field.
@@
$\begin{split}F_{12}\approx&\frac{\hat{h}^{4}\hat{\tau}_{gx}^{2}+24\left(\hat{h}^{2}+2\hat{h}\hat{q}_{x}+6\hat{q}_{x}^{2}\right)}{120\hat{h}}\\ &+\frac{6\hat{h}^{2}\hat{\tau}_{gx}(\hat{h}+\hat{q}_{x})}{120\hat{h}}\end{split}$ (59a)
[…]
$\begin{split}\hat{\tau}_{w,z}\approx&\left(\frac{\hat{\tau}_{g,x}}{2}-\frac{3\hat{q}_{z}}{\hat{h}^{2}}\right)\,.\end{split}$ (59e)
@@ VII.2 Leading-Oder convective velocity
As for the cases treated in subsection VII.1, the leading-order term is consistent with the case without a magnetic field [22].
@@
Figure 8 illustrates the phase speed $c$ (61) as a function of the non-dimensional thickness $\hat{h}$ and the Hartmann number Ha. For small values of Ha, $c$ has a quadratic dependence on $\hat{h}$, with waves propagating upstream ($c<0$) for $\hat{h}<1$ and downstream ($c>0$) for $\hat{h}>1$. For $\mbox{Ha}\geq 2$, $c$ reaches a plateau at approximately $c\approx 0.1$ for $\hat{h}>0.3$. As we have already seen in figure 6(b), the magnetic effects for these conditions counterbalance the gravitational effects, resulting in an equilibrium solution where waves always propagate in the direction of substrate movement. This highlights how the magnetic force may affect the absolute/convective instability properties of these flat-film solutions relative to the case without a magnetic field [40].
@@ VII.3 Undulation control using Reinforcement Learning
This section presents the results of the undulation control problem introduced in subsection VI.2, using the PPO algorithm described in subsection VI.1. Three scenarios are considered: two with a single actuator (either a gas jet or an electromagnetic actuator) and one with both actuators operating simultaneously. The performance of the trained agent is assessed by analysing the evolution of reward during training episodes (learning curves) and by evaluating the agent’s performance over 10 test episodes at the end of training. The results of this evaluation are reported as the mean and standard deviation of episode rewards, which are compared with those from the uncontrolled case. To provide further insight into the control mechanism, we also present the evolution of the control function and snapshots of the liquid film along the centreline during one evaluation episode.
VII.3.1 Control with single gas-jets
The first case involves a single gas-jet actuator with a prescribed harmonic control function, and the agent is free to take any action at any time.
Figure 9: Learning curves for the harmonic (green line with squares) and non-harmonic (blue line with circles) control functions with a gas jet.
@@ VII.3.1 Control with single gas-jets
Table 2 summarises the mean and standard deviation of the evaluation episodes for controlled and uncontrolled cases. As in the learning phase, both control functions yield a higher mean reward than in the uncontrolled case. Furthermore, in the controlled case, the reward variability around the mean is reduced, as indicated by a smaller standard deviation. This shows that the PPO identified an effective control action that reduced undulation amplitude in the reward area.
Focusing on the harmonic case, the evolution of the optimal control function during an evaluation episode is reported in Figure 10. Figure 10(a) shows the amplitude (solid red line with squares), the nondimensional frequency (blue dashed line with triangles), and the phase shift (dotted dashed line with circles) of the harmonic control function during an evaluation episode. These three parameters remain nearly constant, with an amplitude of $\acute{A}\approx 0.68$, a frequency of $\acute{f}\approx 0.065$, and a phase shift of $\acute{\phi}\approx\pi$. The frequency aligns with the average frequency of the undulation waves, which ranges between $0.05$ and $0.08$. This shows how a single harmonic function is sufficient to significantly reduce the amplitude of the undulation instability.
@@
To gain a better understanding of the control mechanism, we examine the evolution of the control function alongside observations of the liquid film collected at one of the three observation locations. Figure 10(b) shows the evolution of the nozzle exit velocity (solid red line) and the observation at obs2 (dashed-green line). The positions of the crests and valleys of the two functions are very similar, with only a slight phase shift due to the observation point being located ahead of the control actuator. This shows that the jet primarily flattens the crests with only a minor effect on the valleys. However, because the nozzle exit velocity is always nonzero, this indicates that the optimal control law involves thinning the liquid film and a harmonic component nearly in phase with the instability wave.
Figure 12: Evolution of (a, b and c) the controlled 3D liquid film at different time steps and (d,e and f) the corresponding film profiles along the streamwise direction $x$ at $z=1\,\text{cm}$ for the controlled (solid blue line) and uncontrolled (dashed green line) cases using a single gas jet actuator.
Moving to the case where the PPO is free to take any actions, Figure 11 shows the evolution of (a) the control actions and the relation between the control actions and (b) the obs1 with the fitted line (red curve) for the non-harmonic control function. The control action is proportional to the mean liquid-film thickness at the impingement point, with a small phase lag.
@@
VII.3.2 Control with single electromagnets
Turning to the control with a single electromagnetic actuator, Figure 13 shows the learning curve (continuous blue line with circles and continuous green line with squares), the mean value (red dashed line) and the standard deviation (yellow shaded area) of the reward in the absence of control. The case with a prescribed harmonic control function fails to produce good results, with the average reward well below that of the uncontrolled case, even falling outside the uncertainty region of the no-control case. This suggests that the agent was unable to find an efficient control law and therefore continued to test random action combinations, resulting in increased undulation amplitude in the reward area. In contrast, the agent discovered an optimal non-harmonic control law without a prescribed control function shape at the 40th learning episode. This episode marks the point at which the mean reward exceeded that of the uncontrolled case. Furthermore, the mean value approaches the upper bound of the uncertainty region in the uncontrolled case, underscoring the algorithm’s robustness.
@@ VII.3.2 Control with single electromagnets
A similar behaviour is observed when examining performance during the evaluation episodes. Table 3 shows the mean and standard deviation of the liquid film with and without control, using both harmonic and non-harmonic control functions with the 2D electromagnetic actuator. The harmonic control function yielded a mean reward 14% lower than in the uncontrolled case, with a larger standard deviation. In contrast, the non-harmonic control function yielded a reward 8% higher than the uncontrolled case, with a smaller standard deviation.
@@
The control strategy obtained with the electromagnets differs significantly from that obtained with a single 2D gas jet. Figure 14 shows (a) the evolution of the control function (solid red line with squares) and the observations (coloured lines without markers) and (b) the relationship between the control function and obs1 during an evaluation episode. The electromagnetic control exhibits a bang-bang-type behaviour, acting only near the valleys with a maximum intensity of $0.14\,\rm{T}$. The actions are inversely proportional to obs1 when they exceed the flat state of the target, $\bar{h}=0.1\,(34.2\,\mu\rm{m})$.
Figure 14: Evolution of the 3D undulation control using 2D electromagnets: (a) observations (coloured curves without markers) and magnetic field intensity (red solid curve with square markers); (b) relationship between the magnetic field intensity and the second observation (black dots, with shading indicating simulation time from light at the beginning to dark at the end), together with the fitted curve (red solid line).
@@
The optimal control law mechanism arises from the resistive effects of the Lorentz force on the liquid film. The Lorenz forces push the liquid in the positive $x$ direction, creating a positive/negative gradient in flow rate across the electromagnet’s centerline. This results in an increase/decrease in the liquid film thickness upstream and a decrease downstream. This behaviour is evident in the comparison of wave evolution between the controlled and uncontrolled cases. Figure 15 shows the $x$-profile of the liquid film with (solid blue line) and without (dashed green line) control. The action of the electromagnet raises and deforms the valleys, which reduces the amplitude of the nearby peaks.
This control mechanism is constrained by the mismatch between the characteristic time scale of the film, $t_{\mathrm{ref}}$, and that of the electromagnetic effects, $t_{\mathrm{ref,M}}$, quantified by a small Stuart number ($\mbox{N}=1.9\times 10^{-2}$). For the magnetic field to exert a significant influence, its characteristic time scale should be comparable to that of the film, i.e., $\mbox{N}=O(1)$. However, given the Reynolds and Capillary numbers considered in this study, this condition would require a very large Hartmann number, expressed as:
$\mbox{Ha}=\sqrt{\mbox{Re}}\,\mbox{Ca}^{1/6}=16,$ (63)
which presents a major challenge for industrial implementation due to the high cost, strong thermal gradients induced by the Joule effect, and the potential destabilising impact on the strip.
@@
Figure 15: Evolution of the controlled (blue continuous line) and the uncontrolled (green dashed line) along $x$ at $z=1\rm{cm}$ for the control of the 3D undulation with a 2D electromagnetic actuator.
VII.3.3 Control with gas jets and electromagnets
@@
In the previous test cases, we found that optimal control actions for the gas jet and electromagnetic actuators are complementary. The gas jet pushes the crests, whereas the electromagnets raise the valleys. In this case, we test the control of undulations with both actuators operating in tandem. Additionally, to simplify the control problem, the agent has access to only three observations of liquid film thickness, rather than six.
@@
Figure 16 shows the learning curve along with the mean (red dashed line) and standard deviation (yellow shaded area) of the reward obtained without control. The PPO algorithm finds an optimal control law that outperforms the uncontrolled case after 100 training episodes. Furthermore, as observed in the single electromagnet case, the learning curve approaches the upper bound of the uncertainty region, demonstrating the robustness of the discovered control action. Similar results are evident in terms of evaluation episodes.
@@
Table 4 shows the mean and standard deviation of the reward obtained during the evaluation episodes, both with and without control. The optimal control law yields a mean reward 13% higher than in the uncontrolled case, and a 20% reduction in the standard deviation. This indicates that, even in the worst-case evaluation test, the agent outperforms the uncontrolled case on average.
Moving to the evolution of optimal control functions during an evaluation episode, Figure 17 shows (a) the actions of the 2D gas jet (blue solid line) and obs3 (green dashed line) and (b) the actions of the 2D electromagnet (red solid line) and obs2 (green dashed line). As in the case of a single gas jet, the optimal control law is nearly proportional to the liquid film observation, with a small phase shift due to the distance between the observation point and the actuator. In contrast, the magnetic control action differs significantly from that in the standalone case. Here, the magnetic field is never zero and varies nonproportionally relative to the observation, unlike the gas jet’s upstream action, which alters the free-surface topography and forces the electromagnet to adopt a different control strategy. Additionally, since both actuators are always nonzero, the optimal control action implies further thinning of the liquid film.
The control mechanism, which combines pushing crests and raising valleys, is most apparent when examining the evolution of the liquid film along a line in the streamwise direction. Figure 18 shows the evolution of the liquid film with control (solid blue line) and without control (dashed green line) along $x$ at $z=1\,[\rm{cm}]$. For a single actuator, the gas jets depress the crests through additional wiping. However, this is counteracted by the electromagnet, resulting in waves with a smaller peak-to-valley amplitude in the controlled case than in the uncontrolled case.